What is the magnitude of the electric field 23.4 cm directly above an isolated 3.50×10−5 C charge?
Answer #1

Electric field of a charge Q at a distance r from it is given by

E = kQ/r2

E = (9*109) * (3.50*10-5C) / (23.4*10-2m)2

E = 5.753*106 N/C